OpenStudy (anonymous):

Simplify ____ √-32 A.)4i√2 B.)-6i√5 C.)-6√5 D.)6√5

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[\sqrt{-32} = \sqrt{-1*2*16} = \sqrt{-1} * \sqrt{2} * \sqrt{16}\] \[\sqrt{-1} = i , \sqrt{16} = 4 \] \[4i\sqrt{2}\]